# Scalable Computations for Generalized Mixed Effects Models with Crossed Random Effects Using Krylov Subspace Methods

### v1

@@ 6 Real-world applications
In the following, we conduct experiments on five real-world data sets with high-dimensional crossed random effects analyzing the runtime and accuracy for parameter estimation when using GLMMs. We compare the proposed Krylov subspace methods with Cholesky-based calculations from the GPBoost, lme4, and glmmTMB packages, using the same model and computational settings as described in Section 5.1. We use the ZIC preconditioner for the regression data sets and the SSOR preconditioner for the classification data sets.
@@ 6.1 Data sets
We consider three regression and two classification data sets. Table 1 gives an overview of the data sets reporting the sample size, the number of predictor variables, the categorical variables modeled using random effects including the number of group levels $m_{j}$, and the total number of non-zero entries in $Z^{T}Z$. In Figure 13 in Appendix A.10, we additionally visualize the non-zero entries of the matrix $Z^{T}Z$ for the different data sets. Note that the black areas in the plots should not hide the fact that the high-dimensional matrices are sparse, which can be seen when the number of nonzero entries is set in relation to the matrix dimensions. The data sets ‘upselling’ (‘KDDCup09_upselling’) and ‘employee_access’ (‘Amazon_employee_access’) have previously been used by Pargent et al. (2022). For more information on the ‘instEval’ and ‘cars’ data sets, we refer to Simchoni and Rosset (2023), and the data set ‘building_permits’ (‘chicago_building_permits’) has been used by Reyes (2019). In general, an extra category is added for missing values in categorical variables and numeric predictor variables are simply imputed using the mean.
data set $n$ $p$ $K$ $y$-class Cat. var. $m_{j}$ $\text{nzz}(Z^{T}Z)$
Regression cars 97’729 66 2 log(price) model_id 15’226 185’807
location_id 12’235
building_permits 527’168 6 3 log(cost) contact_name 98’362 2’544’893
latitude 206’197
longitude 206’170
instEval 73’421 22 2 teacher student 2’972 150’942
rating teacher 1’128
Classification employee_access 32’769 0 9 approval 0.94 resource 7’518 469’262
manager 4’243
@@
upselling 50’000 34 4 up-selling 0.07 Var216 2’016 454’852
Var217 13’991
Var198 4’291
Var199 5’074
Table 1: Summary of real-world data sets. $n$ is the number of samples, $p$ is the number of predictor variables, $K$ is the number of categorical variables modeled with random effects, $y$ describes the response variable, ‘%-class’ is the frequency of the ‘1’ response variables for the classification data sets, ‘Cat. var.’ describes the categorical variables modeled with random effects, $m_{j}$ is the number of group levels, and $\text{nzz}(Z^{T}Z)$ is the number of non-zero entries in the matrix $Z^{T}Z$.
@@ 6.2 Results
Figure 7 shows the estimation runtimes for the different data sets. In Appendix A.10, we additionally report the wall-clock time, the log-marginal likelihood at the optimum, and the estimated parameters. Parameter estimation crashes when using the lme4 and glmmTMB packages on the ‘upselling’ data set. We have additionally tried modeling only the two variables Var216 and Var217 as random effects but still observed crashes on the ‘upselling’ data set. For the ‘employee_access’ data set, we have also tried modeling the low-cardinality categorical variables such as ‘role_family’ and ‘role_cat1’ as dummy-coded fixed effects instead of random effects, but glmmTMB and lme4 crash when doing this. In Figure 8, we report the average differences in estimation runtime relative to the fastest model over the four data sets for which glmmTMB and lme4 do not crash.
@@
The results show that Krylov subspace methods clearly have the fastest runtime on all real-world data sets. Compared to Cholesky-based calculations using the GPBoost package, the average speed-up with Krylov subspace methods is more than a factor of $20$. On the high-dimensional ‘building_permits’ data set, parameter estimation with Krylov subspace methods is over 57 times faster compared to Cholesky-based calculations in GPBoost. Moreover, estimation is, on average, more than $30$ and $80$ times slower when using lme4 and glmmTMB, respectively, compared to Krylov subspace-based methods. On average, the estimation time of lme4 is the second smallest for the three regression data sets. However, as observed in experiments with simulated data, lme4 is very slow on classification data and can crash, e.g., on the ‘upselling’ data set. glmmTMB has overall the longest runtime. Concerning the parameter estimates reported in Tables 8 and 7 in Appendix A.10, we find that all methods and packages yield essentially the same estimates except for lme4 on the ‘building_permits’ and ‘employee_access’ data sets. In summary, our novel Krylov subspace-based methods are substantially faster than Cholesky-based calculations and numerically more stable than existing software libraries for GLMMs.
Figure 8: Average relative difference in runtime to the fastest model over all real-world data sets except for the data set ‘upselling’ where lme4 and glmmTMB crash.
@@ 7 Conclusion
We present Krylov subspace methods for mixed effects models with high-dimensional crossed random effects. We analyze several preconditioners, derive novel theoretical results, and introduce different methods for accurately approximating predictive variances. In experiments, we find that our methods are faster and more stable than Cholesky-based calculations and other state-of-the-art software implementations for GLMMs. In addition, we provide theoretical and empirical results showing that the symmetric successive over-relaxation (SSOR) preconditioner is superior compared to the popular diagonal preconditioner for both the CG and the SLQ methods. Potential directions for future research include analyzing alternative approximations for log-determinants, a systematic comparison under which situations the ZIC or the SSOR preconditioners are preferred, and the development of novel preconditioners.
@@ A.1 Proofs of results in Section 4
We denote by $\sigma_{\min}(B)=\sigma_{\min(n_{1},n_{2})}(B)\leq\dots\leq\sigma_{1}(B)=\sigma_{\max}(B)$ the singular values of a matrix $B\in\mathbb{R}^{n_{1}\times n_{2}}$, and by $\lambda_{\min}(A)=\lambda_{m}(A)\leq\dots\leq\lambda_{1}(A)=\lambda_{\max}(A)$ the eigenvalues of a symmetric matrix $A\in\mathbb{R}^{m\times m}$. Further, we denote by $D=\Sigma^{-1}+\text{diag}(Z^{T}WZ)\in\mathbb{R}^{m\times m}$, $\Delta=\text{diag}(Z^{T}WZ)\in\mathbb{R}^{m\times m}$, and $\Lambda=Z^{T}WZ-\text{diag}(Z^{T}WZ)\in\mathbb{R}^{m\times m}$.
We will use the following two lemmas for proofing Theorem 4.2.
@@
{±$\{\pm\sqrt{\lambda_{i}(B^{T}B)};i=1,\dots,\min(n_{1},n_{2})\}=\{\pm\sigma_{i}(B);i=1,\dots,\min(n_{1},n_{2})\}.$
@@
$L=\begin{pmatrix}0&0\\ Z_{2}^{T}WZ_{1}&0\end{pmatrix},\leavevmode\nobreak\ \leavevmode\nobreak\ D=\begin{pmatrix}D_{1}&0\\ 0&D_{2}\end{pmatrix},\leavevmode\nobreak\ \leavevmode\nobreak\ \Sigma^{-1}+Z^{T}WZ=\begin{pmatrix}D_{1}&Z_{1}^{T}WZ_{2}\\ Z_{2}^{T}WZ_{1}&D_{2}\end{pmatrix}.$
@@
and (21) thus follows since
$\begin{split}P_{\text{SSOR}}^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P_{\text{SSOR}}^{-T/2}&=D^{1/2}(D+L)^{-1}(\Sigma^{-1}+Z^{T}WZ)(D+L)^{-T}D^{1/2}\\ &=I_{m}-\begin{pmatrix}0&0\\ 0&D_{2}^{-1/2}Z_{2}^{T}WZ_{1}D_{1}^{-1}Z_{1}^{T}WZ_{2}D_{2}^{-1/2}\end{pmatrix}.\end{split}$ (43)
@@
Note that $\Delta^{-1/2}\Lambda\Delta^{-1/2}$ is a normalized adjacency matrix of a bipartite weighted graph. Its largest eigenvalues is thus $1$ [Chung, 1997]. By Lemma A.2, or Ostrowski’s theorem [Horn and Johnson, 2012], it follows that
@@
$\min((D_{1}^{-1})_{ii})\min((D_{2}^{-1})_{ii})\min(\Delta_{ii}^{2})\leq\lambda_{\max}(D_{2}^{-1/2}Z_{2}^{T}WZ_{1}D_{1}^{-1}Z_{1}^{T}WZ_{2}D_{2}^{-1/2})\leq\max((D_{1}^{-1})_{ii})\max((D_{2}^{-1})_{ii})\max(\Delta_{ii}^{2})$
@@
Proof of Theorem 4.3.
Similarly as in the proof of Theorem 4.2, we can apply Lemma A.1 and obtain
where we have used
Applying Theorem 3.2 of Brito et al. [2022] to $\lambda_{2}\begin{pmatrix}0&Z_{1}^{T}Z_{2}\\ Z_{2}^{T}Z_{1}&0\end{pmatrix}$ then shows that
asymptotically almost surely with $\epsilon_{m}^{\prime}\rightarrow 0$ as $m\rightarrow\infty$. Using (43) and (44) proofs (27). (28) follows directly from (27).
Proof of Theorem 4.4.
@@
$v_{k}=D^{1/2}(\frac{1}{k}1_{m_{1}}^{T},\dots,\frac{1}{k}1_{m_{k}}^{T},-1_{m_{k+1}}^{T},0,\dots,0)^{T}\in\mathbb{R}^{m},\leavevmode\nobreak\ \leavevmode\nobreak\ k=1,\dots,K-1,$
where $1_{m_{k}}$ denotes a vectors of $1$’s of length $m_{k}$. We have $D^{-1/2}Z^{T}WZD^{-1/2}v_{k}=0$, and the vectors $v_{k}$ are linearly independent. Since $D^{-1/2}Z^{T}WZD^{-1/2}$ is positive semi-definite, this shows that the $K-1$ smallest eigenvalues of $D^{-1/2}Z^{T}WZD^{-1/2}$ are zero. Applying Weyl’s inequality [Horn and Johnson, 2012] to $P_{\text{Diag}}^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P_{\text{Diag}}^{-T/2}=D^{-1/2}\Sigma^{-1}D^{-1/2}+D^{-1/2}Z^{T}WZD^{-1/2}$ gives
@@
for $k=1,\dots,K-1$, which proofs (30).
Similarly as in the proof of Theorem 4.2, we can show, for general $K$, that
@@
proofs the statement in (29). ∎
Proof of Theorem 4.5.
@@
Using $\lambda_{\max}(\Delta^{-1/2}\Lambda\Delta^{-1/2})=K-1$, see the proof of Theorem 4.4, then gives
@@
We have
Applying Weyl’s inequality [Horn and Johnson, 2012], Ostrowski’s theorem [Horn and Johnson, 2012], and the fact that the largest eigenvalue of $\Delta^{-1/2}\Lambda\Delta^{-1/2}$ is $K-1$, see the proof of Theorem 4.4, we obtain (35). (36) is shown analogously as (30) in the proof of Theorem 4.4. ∎
@@
First, (37) follows directly from Corollary 4.1 and Theorem 4.5. Next, by Corollary 4.1, we have
$\lambda_{\min}^{\text{SSOR}}=1-\frac{\sigma^{2}_{1}d}{\sigma^{2}+\sigma^{2}_{1}d}\frac{\sigma^{2}_{2}d}{\sigma^{2}+\sigma^{2}_{2}d}=\frac{(\sigma^{2})^{2}+\sigma^{2}(\sigma^{2}_{1}+\sigma^{2}_{2})d}{(\sigma^{2})^{2}+\sigma^{2}(\sigma^{2}_{1}+\sigma^{2}_{2})d+\sigma^{2}_{1}\sigma^{2}_{2}d^{2}}$
and thus
$\frac{\lambda_{\max}^{\text{SSOR}}}{\lambda_{\min}^{\text{SSOR}}}=\frac{(\sigma^{2})^{2}/d+\sigma^{2}(\sigma^{2}_{1}+\sigma^{2}_{2})+\sigma^{2}_{1}\sigma^{2}_{2}d}{(\sigma^{2})^{2}/d+\sigma^{2}(\sigma^{2}_{1}+\sigma^{2}_{2})},$
and (38) follows. By Theorem 4.5, we have
$\frac{\lambda_{\max}^{\text{Diag}}}{\lambda_{\min}^{\text{Diag}}}=\frac{\left(1+\frac{1}{\sqrt{\left(\frac{\sigma^{2}}{\sigma_{1}^{2}d}+1\right)\left(\frac{\sigma^{2}}{\sigma_{2}^{2}d}+1\right)}}\right)}{\left(1-\frac{1}{\sqrt{\left(\frac{\sigma^{2}}{\sigma_{1}^{2}d}+1\right)\left(\frac{\sigma^{2}}{\sigma_{2}^{2}d}+1\right)}}\right)}=\frac{\sqrt{\left(\frac{\sigma^{2}}{\sigma_{1}^{2}d}+1\right)\left(\frac{\sigma^{2}}{\sigma_{2}^{2}d}+1\right)}+1}{\sqrt{\left(\frac{\sigma^{2}}{\sigma_{1}^{2}d}+1\right)\left(\frac{\sigma^{2}}{\sigma_{2}^{2}d}+1\right)}-1}.$ (49)
Using
one can show that $\frac{\lambda_{\max}^{\text{Diag}}}{\lambda_{\min}^{\text{Diag}}}$ has the linear asymptote in $d$ given in (39). (40) follows from Theorem 4.6.
A.2 Effective condition number for the diagonal preconditioner and proof of Theorem 4.8
@@
Analogously as in the proof of Theorem 4.3, applying Theorem 3.2 of Brito et al. [2022] to $\lambda_{2}\begin{pmatrix}0&Z_{1}^{T}Z_{2}\\ Z_{2}^{T}Z_{1}&0\end{pmatrix}$ then shows that
@@
Proof of Theorem 4.8.
Theorem 4.3 shows that

### v2

@@ 6 Real-world applications
In this section, we conduct experiments on six real-world data sets with crossed random effects, analyzing the runtime and accuracy for parameter estimation. We compare the proposed Krylov subspace methods with Cholesky-based calculations implemented in the GPBoost, lme4, and glmmTMB packages, using the same model and computational settings as described in Section 5.1.
@@ 6.1 Data sets
We consider four regression (‘cars’, ‘building_permits’, ‘instEval’, ‘MovieLens_32m’) and two classification (‘employee_access’, ‘upselling’) data sets. Table 1 gives an overview of the data sets reporting the sample size, the number of covariates included in a linear predictor, the categorical variables modeled using random effects including the number of group levels $m_{k}$, and the total number of non-zero entries in $Z^{T}Z$. In Figure 14 in Appendix A.10, we additionally visualize the non-zero entries of the matrix $Z^{T}Z$ for the different data sets. For the ‘MovieLens_32m’ data set (harper2015movielens), we use linear and quadratic time trends as predictor variables. The data sets ‘upselling’ (‘KDDCup09_upselling’) and ‘employee_access’ (‘Amazon_employee_access’) have previously been used by pargent2022regularized. For more information on the ‘instEval’ and ‘cars’ data sets, we refer to simchoni2023integrating, and the data set ‘building_permits’ (‘chicago_building_permits’) has been used by reyes2019statistical. In general, an extra category is added for missing values in categorical variables and numeric predictor variables are simply imputed using the mean. We use the ZIC preconditioner for the ‘cars’, ‘building_permits’, and ‘instEval’ data sets and the SSOR preconditioner for the other data sets.
data set $n$ $p$ $K$ $y$ % 1 Cat. var. $m_{k}$ $\text{nnz}(Z^{T}Z)$
Regression
cars 97,729 66 2 log(price) model_id 15,226 185,807
location_id 12,235
building_permits 527,168 6 3 log(cost) contact_name 98,362 2,544,893
latitude 206,197
longitude 206,170
instEval 73,421 22 2 rating student 2,972 150,942
teacher 1,128
MovieLens_32m 32,000,204 2 2 rating userId 200,948 64,285,788
movieId 84,432
Classification
employee_access 32,769 0 9 approval 0.94 resource 7,518 469,262
manager 4,243
@@
upselling 50,000 34 4 up-selling 0.07 Var216 2,016 454,852
Var217 13,991
Var198 4,291
Var199 5,074
Table 1: Summary of real-world data sets. $n$ is the number of samples, $p$ is the number of predictor variables, $K$ is the number of categorical variables modeled with random effects, $y$ describes the response variable, ‘% 1’ is the frequency of the ’s in the response variable for the classification data sets, ‘Cat. var.’ describes the categorical variables modeled with random effects, $m_{k}$ is the number of group levels, and $\text{nnz}(Z^{T}Z)$ is the number of non-zero entries in the matrix $Z^{T}Z$.
@@ 6.2 Results
Figure 6 shows the estimation runtimes for the different data sets. In Appendix A.10, we additionally report the wall-clock time, the log-marginal likelihood at the optimum, and the estimated parameters in tabular form. Parameter estimation crashes when using the lme4 and glmmTMB packages on the ‘upselling’ data set. We have additionally tried modeling only the two variables Var216 and Var217 as random effects but still observed crashes on the ‘upselling’ data set. For the ‘employee_access’ data set, we have also tried modeling the low-cardinality categorical variables such as ‘role_family’ and ‘role_cat1’ as dummy-coded fixed effects instead of random effects, but glmmTMB and lme4 crash when doing this. For the ‘MovieLens_32m’ data set, glmmTMB and Cholesky-based computations in GPBoost crash due to memory errors. In Figure 7, we report the average differences in estimation runtime relative to the fastest method. For every method, the average is calculated over the data sets for which the method does not crash.
@@
The results show that Krylov subspace methods clearly have the fastest runtime on all real-world data sets. Compared with Cholesky-based calculations using the GPBoost package, the average speedup of the Krylov subspace methods exceeds a factor of $25$ on those data sets for which Cholesky-based calculations are feasible. On the relatively high-dimensional ‘building_permits’ data set, parameter estimation with Krylov subspace methods is approximately $60$ times faster compared to Cholesky-based calculations. On average, estimation with lme4 and glmmTMB is approximately $220$ and $85$ times slower, respectively, than with the Krylov subspace-based methods. As observed in experiments with simulated data, lme4 is slow on classification data and can crash, e.g., on the ‘upselling’ data set. Furthermore, estimation on the ‘MovieLens_32m’ data set takes approximately $1,000$ times longer with lme4 than with Krylov subspace-based methods. Concerning the parameter estimates reported in Tables 8 and 7 in Appendix A.10, we find that all methods and packages yield essentially the same estimates except for lme4 on the ‘building_permits’ and ‘employee_access’ data sets. In summary, our novel Krylov subspace-based methods are substantially faster than Cholesky-based calculations and numerically more stable.
Figure 7: Average relative difference in runtime to the fastest model for the real-world data sets.
@@ 7 Conclusion
We present Krylov subspace methods for mixed effects models with high-dimensional crossed random effects. We derive novel theoretical results on the convergence and accuracy of these methods, and we analyze them empirically. In experiments, we find that Krylov subspace-based methods are much faster and more stable than Cholesky-based calculations and other state-of-the-art software implementations for GLMMs. For instance, we are able to estimate a model with two crossed random effects and three linear regression coefficients on the large-scale MovieLens_32m data set in less than five minutes, whereas lme4 requires more than three days and Cholesky-based computations in glmmTMB and GPBoost fail due to memory limitations.
We have used the Laplace approximation for non-Gaussian likelihoods following a common practice in widely used software implementations such as lme4 and glmmTMB. A potential caveat of this is that, depending on the asymptotic setting for mixed effects models, the Laplace approximation can be biased (shun1995laplace). Directions for future research include extending our methods to other marginal likelihood approximations, analyzing alternative approximations for log-determinants, a systematic comparison under which situations the ZIC or the SSOR preconditioners are preferred, and the development of novel preconditioners.
@@ A.1 Proofs of results in Section 4
Notation.
We denote by $\sigma_{\min}(B)=\sigma_{\min(n_{1},n_{2})}(B)\leq\dots\leq\sigma_{1}(B)=\sigma_{\max}(B)$ the singular values of a matrix $B\in\mathbb{R}^{n_{1}\times n_{2}}$, and by $\lambda_{\min}(A)=\lambda_{m}(A)\leq\dots\leq\lambda_{1}(A)=\lambda_{\max}(A)$ the eigenvalues of a symmetric matrix $A\in\mathbb{R}^{m\times m}$. Further, let
$D=\Sigma^{-1}+\text{diag}(Z^{T}WZ)\in\mathbb{R}^{m\times m},$
$\Delta=\text{diag}(Z^{T}WZ)\in\mathbb{R}^{m\times m},$
and
$\Lambda=Z^{T}WZ-\text{diag}(Z^{T}WZ)\in\mathbb{R}^{m\times m}.$
We will use the following two lemmas for proving Lemma 4.1.
@@
$\{\pm\sqrt{\lambda_{i}(B^{T}B)};i=1,\dots,\min(n_{1},n_{2})\}=\{\pm\sigma_{i}(B);i=1,\dots,\min(n_{1},n_{2})\}.$
@@
$L=\begin{pmatrix}0&0\\ Z_{2}^{T}WZ_{1}&0\end{pmatrix},\penalty 10000\ \penalty 10000\ D=\begin{pmatrix}D_{1}&0\\ 0&D_{2}\end{pmatrix},\penalty 10000\ \penalty 10000\ \Sigma^{-1}+Z^{T}WZ=\begin{pmatrix}D_{1}&Z_{1}^{T}WZ_{2}\\ Z_{2}^{T}WZ_{1}&D_{2}\end{pmatrix}$.
@@
and we thus obtain
[…]
@@
Note that $\Delta^{-1/2}\Lambda\Delta^{-1/2}$ is a normalized adjacency matrix of a bipartite weighted graph. Its largest eigenvalues is thus 11 [chung1997spectral]. By Lemma A.2, or Ostrowski’s theorem [horn2012matrix], it follows that
@@
$\min((D_{1})_{ii}^{-1})\min((D_{2})_{ii}^{-1})\min(\Delta_{ii}^{2})\leq\lambda_{\max}(D_{2}^{-1/2}Z_{2}^{T}WZ_{1}D_{1}^{-1}Z_{1}^{T}WZ_{2}D_{2}^{-1/2})\leq\max((D_{1})_{ii}^{-1})\max((D_{2})_{ii}^{-1})\max(\Delta_{ii}^{2})$
@@
Proof of Lemma 4.2.
@@
$v_{k}=D^{1/2}(\frac{1}{k}1_{m_{1}}^{T},\dots,\frac{1}{k}1_{m_{k}}^{T},-1_{m_{k+1}}^{T},0,\dots,0)^{T}\in\mathbb{R}^{m},\penalty 10000\ \penalty 10000\ k=1,\dots,K-1,$
where $1_{m_{k}}$ denotes a vectors of $1$’s of length $m_{k}$. We have $D^{-1/2}Z^{T}WZD^{-1/2}v_{k}=0$, and the vectors $v_{k}$ are linearly independent. Since $D^{-1/2}Z^{T}WZD^{-1/2}$ is positive semi-definite, this shows that the K1K-1 smallest eigenvalues of $D^{-1/2}Z^{T}WZD^{-1/2}$ are zero. Applying Weyl’s inequality [horn2012matrix] to $P_{\text{Diag}}^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P_{\text{Diag}}^{-T/2}=D^{-1/2}\Sigma^{-1}D^{-1/2}+D^{-1/2}Z^{T}WZD^{-1/2}$ gives
@@
for $k=1,\dots,K-1$, which proves (25).
Similarly as in the proof of Lemma 4.1, we can show, for general $K$, that
@@
proves the statement in (24).
Next, we have
Applying Weyl’s inequality [horn2012matrix], Ostrowski’s theorem [horn2012matrix], and the fact that the largest eigenvalue of $\Delta^{-1/2}\Lambda\Delta^{-1/2}$ is K1K-1, see above, we obtain (26). (27) is shown analogously as (25) above. ∎
For proofing parts of Theorem 4.2, we first derive the following slightly more general result.
Theorem A.1.
If the likelihood is Gaussian, it holds that
$1+\frac{K-1}{\frac{\max(\Sigma^{-1})\sigma^{2}}{d_{\min}}+1}\leq\lambda_{\max}^{\text{Diag}}\leq 1+\frac{K-1}{\frac{\min(\Sigma^{-1})\sigma^{2}}{d_{\max}}+1},$ (44)
$\frac{1}{\frac{\max(\Sigma)}{\sigma^{2}}d_{\max}+1}\leq\lambda_{m+1-k}^{\text{Diag}}\leq\frac{1}{\frac{\min(\Sigma)}{\sigma^{2}}d_{\min}+1},\penalty 10000\ \penalty 10000\ k=1,\dots,K-1,$ (45)
where $d_{\max}=\max(Z^{T}Z)$ and $d_{\min}=\min(Z^{T}Z)$.
If, in addition, $K=2$ and the design of the random effects is balanced in the sense that every random effect occurs $d_{k}$ times for component $k=1,2$, i.e., $d_{k}=\frac{n}{m_{k}}=\sum_{i=1}^{n}(Z_{k})_{ij}$ for all $1\leq j\leq m_{k}$, it holds that
Proof of Theorem A.1.
@@
Using $\lambda_{\max}(\Delta^{-1/2}\Lambda\Delta^{-1/2})=K-1$, see the proof of Lemma 4.2, then gives
@@
First, (22) and
$\min(\Sigma^{-1}_{ii}(Z^{T}WZ)_{ii}^{-1})\geq\min(\Sigma^{-1})_{ii}\min((Z^{T}WZ)_{ii}^{-1})=\frac{\sigma^{2}}{\sigma^{2}_{\max}d_{\max}}$
imply
$1-\left(\frac{\sigma^{2}_{\max}d_{\max}}{\sigma^{2}+\sigma^{2}_{\max}d_{max}}\right)^{2}\leq\lambda_{\min}^{\text{SSOR}}\leq 1-\left(\frac{\sigma^{2}_{\min}d_{\min}}{\sigma^{2}+\sigma^{2}_{\min}d_{\min}}\right)^{2}$
which gives (28). In addition,
$\frac{1}{1-\left(\frac{\sigma^{2}_{\max}d_{\max}}{\sigma^{2}+\sigma^{2}_{\max}d_{\max}}\right)^{2}}=\frac{1}{2}\frac{\sigma^{2}_{\max}d_{\max}}{\sigma^{2}}+\frac{3}{4}+\frac{1}{8\sigma^{2}_{\max}d_{\max}/\sigma^{2}+4}$
yields the asymptotic result in (28).
Furthermore, plugging in $D_{k}=\text{diag}(1/\sigma^{2}_{k}+d_{k}/\sigma^{2})$, $k=1,2$, in (23) gives
$\begin{split}\frac{\lambda_{\max}^{\text{SSOR}}}{\lambda_{\min}^{\text{SSOR}}}&=\frac{1}{1-\left(\frac{1}{\frac{\sigma^{2}}{\sigma^{2}_{1}d}+1}\right)\left(\frac{1}{\frac{\sigma^{2}}{\sigma^{2}_{2}d}+1}\right)}=\frac{1}{\frac{(\sigma^{2})^{2}+\sigma^{2}(\sigma^{2}_{1}+\sigma^{2}_{2})d}{(\sigma^{2})^{2}+\sigma^{2}(\sigma^{2}_{1}+\sigma^{2}_{2})d+\sigma^{2}_{1}\sigma^{2}_{2}d^{2}}}=\frac{(\sigma^{2})^{2}/d+\sigma^{2}(\sigma^{2}_{1}+\sigma^{2}_{2})+\sigma^{2}_{1}\sigma^{2}_{2}d}{(\sigma^{2})^{2}/d+\sigma^{2}(\sigma^{2}_{1}+\sigma^{2}_{2})},\end{split}$
which yields
$\lim_{d\rightarrow\infty}\left(\frac{\lambda_{\max}^{\text{SSOR}}}{\lambda_{\min}^{\text{SSOR}}}-\left(\frac{\sigma^{2}_{1}\sigma^{2}_{2}}{\sigma^{2}(\sigma^{2}_{1}+\sigma^{2}_{2})}d+1\right)\right)=0$
and thus proves the statements in (31).
Next, Theorem A.1 implies
$\left(1+\frac{1}{\frac{\sigma^{2}}{\sigma^{2}_{\min}d_{\min}}+1}\right)\left(\frac{\sigma^{2}_{\min}d_{\min}}{\sigma^{2}}+1\right)\leq\frac{\lambda_{\max}^{\text{Diag}}}{\lambda_{\min}^{\text{Diag}}}\leq\left(1+\frac{1}{\frac{\sigma^{2}}{\sigma^{2}_{\max}d_{\max}}+1}\right)\left(\frac{\sigma^{2}_{\max}d_{\max}}{\sigma^{2}}+1\right),$
which gives (29). Similarly, by Theorem A.1, we have
$\frac{\lambda_{\max}^{\text{Diag}}}{\lambda_{\min}^{\text{Diag}}}=\frac{\left(1+\frac{1}{\sqrt{\left(\frac{\sigma^{2}}{\sigma_{1}^{2}d}+1\right)\left(\frac{\sigma^{2}}{\sigma_{2}^{2}d}+1\right)}}\right)}{\left(1-\frac{1}{\sqrt{\left(\frac{\sigma^{2}}{\sigma_{1}^{2}d}+1\right)\left(\frac{\sigma^{2}}{\sigma_{2}^{2}d}+1\right)}}\right)}=\frac{\sqrt{\left(\frac{\sigma^{2}}{\sigma_{1}^{2}d}+1\right)\left(\frac{\sigma^{2}}{\sigma_{2}^{2}d}+1\right)}+1}{\sqrt{\left(\frac{\sigma^{2}}{\sigma_{1}^{2}d}+1\right)\left(\frac{\sigma^{2}}{\sigma_{2}^{2}d}+1\right)}-1},$
which is the first equality in (32). Using
one can show that $\frac{\lambda_{\max}^{\text{Diag}}}{\lambda_{\min}^{\text{Diag}}}$ has the linear asymptote $\frac{4\sigma^{2}_{1}\sigma^{2}_{2}}{\sigma^{2}(\sigma^{2}_{1}+\sigma^{2}_{2})}$, i.e.,
$\lim_{d\rightarrow\infty}\left(\frac{\lambda_{\max}^{\text{Diag}}}{\lambda_{\min}^{\text{Diag}}}-\left(\frac{4\sigma^{2}_{1}\sigma^{2}_{2}}{\sigma^{2}(\sigma^{2}_{1}+\sigma^{2}_{2})}d+1\right)\right)=0.$
Finally, (30) follows from Lemma 4.2. ∎
@@
We first note that
$\begin{split}\min(\Sigma^{-1}_{ii}(Z^{T}WZ)_{ii}^{-1})&\geq\frac{1}{\max(\Sigma)_{ii}\max((Z^{T}WZ)_{ii})}\\ &\geq\frac{1}{\max(\Sigma)_{ii}\max((Z^{T}Z)_{ii})\max(W)}\\ &\geq\frac{1}{\sigma^{2}_{\max}d_{\max}C},\end{split}$ (49)
since
$\max(W)\leq C,$ (50)
where $C$ equals $0.25$ and $1$ for a logit and a probit link, respectively. This and (22) imply
$1-\left(\frac{\sigma^{2}_{\max}d_{\max}C}{1+\sigma^{2}_{\max}d_{\max}C}\right)^{2}\leq\lambda_{\min}^{\text{SSOR}},$
which gives the inequality in (33). Similarly as in the proof of Theorem 4.2, we have
$\frac{1}{1-\left(\frac{\sigma^{2}_{\max}d_{\max}C}{1+\sigma^{2}_{\max}d_{\max}C}\right)^{2}}=\frac{\sigma^{2}_{\max}C}{2}d_{\max}+\frac{3}{4}+\frac{1}{8\sigma^{2}_{\max}d_{\max}C+4},$
which yields the asymptotic result in (33).
For the diagonal preconditioner, (24) and the inequality in (49) imply
$\lambda_{\max}^{\text{Diag}}\leq 1+\frac{1}{\frac{1}{\sigma^{2}_{\max}d_{\max}C}+1}=\frac{2\sigma^{2}_{\max}d_{\max}C+1}{\sigma^{2}_{\max}d_{\max}C+1},$
and (25) and (50) imply
$\frac{1}{1+\sigma^{2}_{\max}d_{\max}C}\leq\lambda_{\min}^{\text{Diag}}.$
This gives (34). For the unpreconditioned matrix, Lemma 4.2 and (50) imply (35). ∎
Proof of Theorem 4.4.
Similarly as in the proof of Lemma 4.1, we can apply Lemma A.1 and obtain
where we have used
Applying Theorem 3.2 of brito2022spectral to $\lambda_{2}\begin{pmatrix}0&Z_{1}^{T}Z_{2}\\ Z_{2}^{T}Z_{1}&0\end{pmatrix}$ then shows that
asymptotically almost surely with $\epsilon_{m}^{\prime}\rightarrow 0$ as $m\rightarrow\infty$. Using (40) and (41) then give
$\begin{split}\lambda_{m-1}^{\text{SSOR}}&\geq 1-\frac{(\sqrt{d_{1}-1}+\sqrt{d_{2}-1})^{2}}{(\sigma^{2}/\sigma_{1}^{2}+d_{1})(\sigma^{2}/\sigma_{2}^{2}+d_{2})}-\epsilon_{m}\\ &\geq 1-\left(\frac{1}{d_{1}}+\frac{1}{d_{2}}+\frac{2}{\sqrt{d_{1}d_{2}}}\right)-\epsilon_{m}.\end{split}$ (52)
Finally, (36) follows directly from (52). ∎
@@
Analogously as in the proof of Theorem 4.4, applying Theorem 3.2 of brito2022spectral to $\lambda_{2}\begin{pmatrix}0&Z_{1}^{T}Z_{2}\\ Z_{2}^{T}Z_{1}&0\end{pmatrix}$ then shows that
@@
Proof of Corollary 4.1.
Theorem 4.4 implies that